# Knowledge Boundary of Large Language Models: A Survey

### v1

@@ C.1 Prompt-Sensitive Known Knowledge
For instruction optimization, APE Zhou et al. (2023b) leverages LLMs to automatically generate and perform Monte Carlo search for the instructions, and evaluate the zero-shot performance of the candidate instructions. ORPO Yang et al. (2024a) utilizes natural language as optimization instructions for LLMs to iteratively optimize the prompts. For demonstration optimization, KATE Liu et al. (2022a) retrieve the K nearest in-context examples by the semantic similarity to the test example, measured by the embedding from an encoder model.
@@
Tree-of-thoughts Yao et al. (2023) improves the linear chain-of-thoughts reasoning into tree structure, each node representing a piece of thoughts, and branches represents alternative thoughts. It allows LLMs to perform various forms of reasoning steps. Progressive-hint-prompting Zheng et al. (2023b) appends the LLM-generated answers to the prompt as hints to iteratively arrive at the correct answers.
@@
Self-refine Madaan et al. (2024) prompts LLMs to generate feedback on its previous answer for iterative answer refinement. Self-verification Weng et al. (2023) transforms the generated answer into abductive reasoning questions to examine the consistency with the given context.
@@ C.2 Model-specific Unknown Knowledge
Joshi et al. (2024) enhances low-resource language adaptation by constructing a synthetic Hindi corpus through English text translation, transliteration, and noise filtering, followed by training on the curated dataset. StructTuning Liu et al. (2024c) constructs structured domain knowledge by automatically extracting knowledge taxonomies from corpora, linking text segments to specific knowledge points for efficient model fine-tuning.

### v2

@@ C.1 Prompt-Sensitive Known Knowledge
For instruction optimization, APE Zhou et al. (2023b) leverages LLMs to automatically forward-generate and perform Monte Carlo search on the prompts, and evaluate the performance of the candidate prompts via reverse generation, which consists of $n$ rounds. For demonstration optimization, KATE Liu et al. (2022a) retrieve the K nearest in-context examples by the semantic similarity to the test example, measured by the embedding from an encoder model.
@@
Chain-of-thoughts Wei et al. (2022b) generates the step-by-step rationales followed by the answer. Tree-of-thoughts Yao et al. (2023) improves the linear chain-of-thoughts reasoning into tree structure, each node representing a piece of thoughts, and branches represents alternative thoughts. It allows LLMs to perform various forms of reasoning steps. Progressive-hint-prompting Zheng et al. (2023b) appends the LLM-generated answers to the prompt as hints to iteratively arrive at the correct answers.
@@
Self-refine Madaan et al. (2024) prompts LLMs to generate feedback on its previous answer for iterative answer refinement. Self-verification Weng et al. (2023) transforms the generated answer into abductive reasoning questions to examine the consistency with the given context. Self-correction Huang et al. (2023a) employs an interative initial CoT, review, and answer improvement process. MAD Du et al. (2024) utilize multiple LLM agents to evaluate other LLMs’ answers and update their own answers until they reach a consensus.
@@ C.2 Model-specific Unknown Knowledge
Fact-based SFT Mecklenburg et al. (2024) constructs a systematically covered fact-level question-answer dataset by extracting key facts from documents and generating diverse training examples, then enhances LLMs through SFT to improve their accuracy and adaptability to out-of-domain knowledge. StructTuning Liu et al. (2025) constructs structured domain knowledge by automatically extracting knowledge taxonomies from corpora, linking text segments to specific knowledge points for efficient model fine-tuning. Factuality alignment methods Lin et al. (2024a); Huang and Chen (2024b); Tian et al. (2024) is also a category of approach under this type, enhancing LLM knowledge via alignment approaches such as DPO.
@@ C.3 Model-agnostic Unknown Knowledge
Type Method Training Cost Inference Cost
Prompt Optimization APE Zhou et al. (2023b) N/A $n$ round (prompt forward generation/monte carlo search + prompt reverse generation)
Prompt-based Reasoning CoT Wei et al. (2022a) N/A step-by-step reasoning + answer
PHP Zheng et al. (2023b) N/A $n$ round * (step-by-step reasoning + answer)
Self-refinement Self-correction Huang et al. (2024b) N/A initial generation + review + revise
MAD Du et al. (2024) N/A $n$ round * $m$ agent
Factuality Decoding DoLA Chuang et al. (2024) N/A initial decoding + contrastive decoding step
ITI Li et al. (2023a) probing the truthful direction perturbed attention decoding step
External Knowledge Retrieval HyDE Gao et al. (2023) N/A hypothetical document generation + retrieval + generation
FLARE Jiang et al. (2023b) N/A $n$ (retrieval + generation)
Parametric Knowledge Editing postEdit Song et al. (2024b) retrieval + SFT generation
MELO Yu et al. (2024) retrieval + SFT generation
Knowledge-enhanced Fine-tuning Fact-based SFT Mecklenburg et al. (2024) fact extraction + SFT generation
StructTuning Liu et al. (2025) structure-aware continual pre-training + SFT generation
Refusal KUQ Amayuelas et al. (2024) SFT generation
R-tuning Zhang et al. (2024a) SFT generation
Asking Clarification Questions ProCoT Deng et al. (2023b) N/A step-by-step reasoning + generation
ACT Chen et al. (2024d) preference data construction + direct preference optimization $n$ generation
Table 2: Cost-effective comparison of representative mitigation techniques.
@@
Appendix D Cost-effective Summarization of Representative Mitigation Techniques
We present a cost-effective comparison of representative mitigation techniques in Section 5, aiming to compare their usefulness and provide recommendations, as summarized in Table 2. This table offers a clearer and more structured comparison of these methods, helping readers better understand their relative strengths and limitations. However, directly and fairly comparing the exact performance of these methods remains challenging due to the current lack of a general and comprehensive benchmark for evaluating different mitigation approaches. Further discussions on this challenge can be found in Section 6.
From the table, we can make the following observations: (1) Prompt optimization, prompt-based reasoning, and self-refinement typically follow two main patterns: step-by-step reasoning and multi-round refinement. These fundamental approaches enhance performance, though their specific design and cost vary depending on the method used. (2) In factuality decoding, DoLA operates purely as a decoding method, whereas ITI includes a probing stage with parameter updates. This distinction can guide the choice between the two methods. (3) The main frameworks of external knowledge retrieval and parametric knowledge editing focus on integrating retrieval and inference while minimizing the cost of both components. (4) The cost of refusal and asking clarification questions methods mainly depends on whether fine-tuning on a constructed dataset is required.